y=3x-4 for x=2,5,6,8
1
Expert's answer
2011-05-05T08:02:10-0400
x= 2 y = 3*2 - 4 = 2
x= 5 y = 3*5 - 4 = 11
x = 6& y = 3*6 - 4 = 14
x = 8& y = 3*8 - 4 = 20